OpenStudy (anonymous):

A wheel made 132 revolutions at 9 7/8 revolutions per minute. How many minutes did it rotate. How many minutes did the wheel rotate?

OpenStudy (anonymous):

132 revolutions / 9 7/8 revolutions per minute = 13.37 minutes

OpenStudy (anonymous):

ok so how do you write that in a fraction

OpenStudy (anonymous):

1056/79

OpenStudy (anonymous):

since 9 7/8= 79/8 132/x=79/8 where x is the minutes

OpenStudy (anonymous):

1056/79. just conversion factors

OpenStudy (anonymous):

im so confuse.

OpenStudy (anonymous):

i know its 13 something

OpenStudy (anonymous):

132 revolutions x \[\frac{ 1 revolution }{ 79/8 minutes}\]

OpenStudy (anonymous):

its 13 29/79

OpenStudy (anonymous):

= answer 1056/79

OpenStudy (anonymous):

it's the same answer

OpenStudy (anonymous):

mine is just in an improper fraction, yours is in a mixed

OpenStudy (anonymous):

sorry, the 1 should say minute and 79/8 revolutions